OpenStudy (anonymous):

Having trouble getting this in (ab+i)/c form. Thanks for any help. (2+6i)/(3-9i)

OpenStudy (zzr0ck3r):

multiply top and bottom by \(3+9i\)

OpenStudy (zzr0ck3r):

\[\dfrac{(2+6i)}{(3-9i)}\dfrac{(3+9i)}{(3+9i)}=\dfrac{6+18i+18i+54i^2}{3+27i-27i-81i^2}\] Remembering \(i^2=-1\) can you now solve the problem?

OpenStudy (anonymous):

I got -48+36i/90 so -16+9i/30 yes? I think I multiplied by 3-9i the first time and that's why I couldn't get it to work out.